A bubble of helium gas has a volume 500 mL near the bottom of a tank where the pressure is 1.84 atm and a temperature of 12 degrees Celsius. Determine the bubble's volume once it has risen to the top where the pressure is 1.01 atm and 16 degrees Celsius.
Answer #1

We can use the combined gas law equation.


PV =k      For same moles of same gas


Where P - pressure,

V - Volume,

T- Temperature

And k - constant

So for two instances of gas having P,V and T variable, above equation will become

PIVI - P2V2 T2

P1 = 1.84 atm (near the bottom)

V1 = 500 mL (near the bottom)

T1 = 12 degrees Celsius = 273.15 +12 = 285.15 K (near the bottom)

P2 = 1.01 atm (risen to the top)

V2 = To find (risen to the top)

T2 = 16 degrees Celsius = 273.15 +16 = 289.15 K (risen to the top)

Putting all these in above equation, we get

(1.84 atm x 500 mL) / 285.15 K = (1.01 atm x V2 mL) / 289.15 K

3.226 = (1.01 x V2) / 289.15

V2 = (3.226 x 289.15)/ 1.01

V2 = 923.56 mL

bubble's volume once it has risen to the top 923.5 mL